A car is travelling with uniform motion with a total frictional resistance of 2.8 × 103 N acting in a direction opposite to the motion of the car. What is the force acting on the car in the direction of motion?[1 K, 1 Inq]
1
Expert's answer
2019-06-17T13:09:10-0400

For travelling with uniform motion


"ma=F-F_f=0"

"F=F_f=2.8 \cdot 10^3 \ N"